4.40 An experiment can result in one of five equally likely simple events, El, E2.. . . , E, Events A, B, and C are defined as follows: A: E, E P(A)-4 C: Es, E P(C)-4 Find the probabilities associated with the following events by listing the simple events in each

Find P(A and B)

E1 is in both of them.
Answer #1

ANSWER:

An experiment can result in one of five equally likely simple events, E1, E2, . . . , E5. Events A, B, and C are defined as follows:

A: E1, E3 B. El. E2. E4, Es C: E3, E4 4 P(A) = .4 P(B) = .8 P(C) = .4 0

Find the probabilities associated with the following events by listing the simple events in each.

a. Ac e. BC h.(ANB) f. AB g.AUBUC

We have,

2 48) ABA​​​​​​​

  1. Consider,

    The event 681-4-44e-i5.png

    681-4-44e-i6.png

    = .2

    P(AnB) (From Conditional probability)

    681-4-44e-i8.png

    = 0.25

    (b)

  2. Consider,

    The event 681-4-44e-i10.png

    5 P(BnC)

    = .2

    P(B|C) P(BIC)- P(C) (From Conditional probability)

    = 4

    = .5

(a)

Consider,

P(A)P(B)= (A)(A)

681-4-46e-i4.png

And

P(AnB) 25

P(AnB) P(A)P(B

Events A and B are not independent, since these probabilities are not the same.

(b)

681-4-46e-i7.png= .25

681-4-46e-i8.png0

Events A and B are not mutually exclusive, since the probability is not equal to zero.

(a) P(AUB)-P(A)+P(B)-P(AnB) (Addition rule)

  1. 681-4-45e-i6.png (Multiplication rule)

    4 +8-(.8 (.25)

    681-4-45e-i8.png

  2. (b) 681-4-45e-i9.png681-4-45e-i10.png (Multiplication rule)

    = (.25)(8)

    = 0.2

  3. 681-4-45e-i13.png (Multiplication rule)

  4. = (5)(4)

    = 0.2

    Yes, I agree, because both the results are same.

a) P(A|B) = P(A and B)/P(B) = 0.2/0.8 = 0.25

b) P(B | C) = P(B and C)/P(C) = 0.2/0.4 = 0.5